ZombieSlayer:

Given f(x) = x − 7 and g(x) = x2 . Find g(f(−1)). g(f(−1)) =

jimthompson5910:

You'll need to post the functions for f(x) and g(x)

ZombieSlayer:

Given f(x) = x − 7 and g(x) = x2 .

ZombieSlayer:

Is that what you meant?

jimthompson5910:

yes

jimthompson5910:

The inner function is f(x) What is f(-1) equal to?

ZombieSlayer:

7?

jimthompson5910:

f(x) = x-7 f(-1) = -1-7 .... x replaced with -1 f(-1) = ??

ZombieSlayer:

-8?

jimthompson5910:

yes

jimthompson5910:

now g( f(-1) ) will turn into g(-8) Since f(-1) = -8, we can replace f(-1) with -8

jimthompson5910:

now you just need to evaluate g(-8)

ZombieSlayer:

64?

jimthompson5910:

correct

ZombieSlayer:

there is a 2nd part

jimthompson5910:

second part?

ZombieSlayer:

Find f(g(−1)). f(g(−1)) =

jimthompson5910:

i see

jimthompson5910:

so you'll need g(-1) first

ZombieSlayer:

the inverse

ZombieSlayer:

I think

jimthompson5910:

it's not the inverse, but I can see what you mean

jimthompson5910:

g(x) = x^2 g(-1) = (-1)^2 ... replace every x with -1 g(-1) = ??

ZombieSlayer:

ok...... would it be 2?

jimthompson5910:

no

jimthompson5910:

we are squaring -1

jimthompson5910:

(-1)^2 = (-1)*(-1) = ??

ZombieSlayer:

oh

ZombieSlayer:

1 = 1? so would it be (-1)^2 = (-1)*(-1) = 1

jimthompson5910:

yes

jimthompson5910:

g(-1) = 1 so f( g(-1) ) becomes f(1)

jimthompson5910:

your last step is to find f(1)

ZombieSlayer:

-6?

jimthompson5910:

correct

ZombieSlayer:

Thank you!

jimthompson5910:

you're welcome

ZombieSlayer:

I'm really struggling lol

jimthompson5910:

you did pretty good in my opinion

ZombieSlayer:

Thank you again

jimthompson5910:

sure thing